# Policy Learning with a Language Bottleneck

### v1

@@ A.1 Maze
A.2 Dataset Details for Birds.
@@
A.3 Full Prompts
@@
SelectSay:
You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Possible ACTIONS an agent can take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION describes the ordered sequence of actions that AGENT 1 picks, and each ACTION describes the ACTION that AGENT 2 picks based on the given OBSERVATION. The examples are separated into HIGH REWARD and LOW REWARD examples.+[samples]+Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
Maze:
You will be given a list of example (OBSERVATION, ACTION) trajectories collected from an AGENT learning to solve a maze. Each trajectory receives a REWARD. Possibles OBSERVATIONS an agent see are: WHITE, RED, BLUE Possible ACTIONS an agent can take are: NORTH, SOUTH, EAST, WEST. The examples are separated into HIGH REWARD and LOW REWARD examples + [samples] + Output a language rule that best summarizes the strategy the AGENT should follow when picking a sequence of ACTIONS to solve the maze and receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
Builder:
There are two agents. The goal of Agent 1 is to provide instructions to Agent 2 that helps Agent 2 to successfully recreate the image. You will be given a list of (ORIGINAL, AGENT 1 INSTRUCTION, REWARD) values where ORIGINAL is the original description of an image, INSTRUCTION is the instruction provided by Agent 1 to Agent 2, and REWARD is the reward Agent 2 receives when trying to re-create the image (higher is better). The examples are separated into HIGH REWARD and LOW REWARD examples. + [samples]+
@@
Birds:
The top row of three images have the following HIGH REWARD descriptions:+high reward samples+The bottom row of three images have the following LOW REWARD descriptions:+low reward samples+Provide a rule I should follow in order to provide image descriptions with HIGH REWARD, not LOW REWARD. Provide the rule after the prefix RULE:"
@@
    SelectSay: $[\mathcal{L}]$+Agent 1 selected [observation]. So I should select
@@
Appendix B Additional Results
@@ B.1 Examples of Generated Rules for All Environments
Table 1: Example $\mathcal{L}$ rules generated for the SelectSay environment, for the Standard setting (both Listener and Speaker agents are RL agents trained from random initialization) and a Fixed Speaker agent.
@@
Table 2: Example $\mathcal{L}$ rules generated for the Maze environment for two different maze sizes.

### v2

@@ A.2 Maze
A.3 Dataset Details for Birds.
@@
A.4 Grasp
Figure 15: Task environment for Grasp, showing new location and distractors for the Generalization setting.
A.5 Full Prompts
@@
  1. 1.
    SaySelect: You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Possible ACTIONS an agent can take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION describes the ordered sequence of actions that AGENT 1 picks, and each ACTION describes the ACTION that AGENT 2 picks based on the given OBSERVATION. The examples are separated into HIGH REWARD and LOW REWARD examples.+[samples]+Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
  2. 2.
    Maze: You will be given a list of example (OBSERVATION, ACTION) trajectories collected from an AGENT learning to solve a maze. Each trajectory receives a REWARD. Possibles OBSERVATIONS an agent see are: WHITE, RED, BLUE Possible ACTIONS an agent can take are: NORTH, SOUTH, EAST, WEST. The examples are separated into HIGH REWARD and LOW REWARD examples + [samples] + Output a language rule that best summarizes the strategy the AGENT should follow when picking a sequence of ACTIONS to solve the maze and receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ’I should’.
  3. 3.
    Builder: There are two agents. The goal of Agent 1 is to provide instructions to Agent 2 that helps Agent 2 to successfully recreate the image. You will be given a list of (ORIGINAL, AGENT 1 INSTRUCTION, REWARD) values where ORIGINAL is the original description of an image, INSTRUCTION is the instruction provided by Agent 1 to Agent 2, and REWARD is the reward Agent 2 receives when trying to re-create the image (higher is better). The examples are separated into HIGH REWARD and LOW REWARD examples. + [samples]+
@@
  4. 4.
    Birds: The top row of three images have the following HIGH REWARD descriptions:+high reward samples+The bottom row of three images have the following LOW REWARD descriptions:+low reward samples+Provide a rule I should follow in order to provide image descriptions with HIGH REWARD, not LOW REWARD. Provide the rule after the prefix RULE:
  5. 5.
    Grasp: The top image shows a grasp keypoint with HIGH REWARD. The bottom image shows a grasp keypoint with LOW REWARD. Based on these images, provide a rule the robot should follow in order to select a grasp keypoint that results in HIGH REWARD, not LOW REWARD. Provide the rule after the prefix RULE:
@@
    SaySelect: $[\mathcal{L}]$+Agent 1 selected [observation]. So I should select
@@
  5. 5.
    Grasp: Provide a keypoint in the image where the robot should grasp the object, following the RULE: $\mathcal{L}$.
Appendix B Complementary Results
@@ B.1 Examples of Generated Rules for All Environments
Table 1: Example $\mathcal{L}$ rules generated for the SaySelect environment, for the Standard setting (both Listener and Speaker agents are RL agents trained from random initialization) and a Fixed Speaker agent.
@@
Table 2: Example $\mathcal{L}$ rules generated for the Maze environment for the 7x7 maze size.
@@
Reward $\mathcal{L}_{10}$ $\mathcal{L}_{30}$ $\mathcal{L}_{50}$
Shoulder Torque Select keypoints that are away from sharp edges or corners of the object to avoid potential damage and improve stability during grasping. Prioritize grasp keypoints that are closer to the center of the object for more stable and precise pickup. Prioritize grasping at the middle of the object’s surface rather than the top edges or handles to ensure a stable grip.
Table 5: Example $\mathcal{L}$ rules generated for the Grasp environment.
B.2 Effect of Model on Generated Rules
We additionally compare the effect of different language models, in particularly those of different sizes, on our results. We did not observe a significantly strong quantitative difference in performance when gen_rule is instantiated with models of different sizes (e.g. llama-2-70b-chat vs. llama-2-13b-chat). However, we did notice interesting qualitative differences across samples that are likely due to the additional fine-tuning step using reinforcmenet learning from human feedback (RLHF, see https://llama.meta.com/llama2/ for more information). We describe these differences per environment below.
  1. 1.
    SaySelect: Because of the simplicity of this environment, it is possible for a rule $\mathcal{L}$ to summarize the full optimal human-interpretable policy as a sequence of if-statements (e.g. “If the current state is ’1’, I should take action 1.”) - we observed that smaller language models (e.g. llama-2-7b-chat and llama-2-13b-chat) always did this, while larger models (e.g. llama-2-70b-chat and mistral-8x-7b-instruct) were better able to generalize and use more efficient language, such as “I should take the same action as the observation”. The smaller models also often included superfluous language, such as “I should always take the action that leads to the highest reward”.
  2. 2.
    Maze: The majority of generated rules captured the underlying color semantics of the maze, enabling generalization. However, smaller model sizes (e.g. llama-2-7b-chat) resulted in more superfluous language (e.g. “I should always prioritize taking actions that lead to the most recent reward, and avoid taking actions that lead to low reward.”), and the generated rules focused more on the first few actions the agent should take, which may not always generalize.
  3. 3.
    Builder: While we do not observe any model-specific differences, there exists variation across samples in the type of formatting and syntax generated rules encourage (e.g. provide coordinates “ using the format (x, y)” vs. “ using the format "x=0.XX, y=0.YY"”), leading to agents converging to different descriptions. Furthermore, some rules encourage list formats in image descriptions (e.g. 1.Draw a green dot at (0.72, 0.21). 2. Draw a green dot at (0.73, 0.72).) while other rules encouraged clustering of identical shapes (e.g. Draw two green dots at (0.72, 0.21) and (0.73, 0.72).)
  4. 4.
    Birds: Rules demonstrated more reward-specificity (i.e. specific to background, color, or species rewards) when generated with larger VLMs (e.g. llava-13b) versus smaller models (e.g. llava-v1.6-vicuna-7b), with the latter primarily proposing rules that encouraged more detailed descriptions (e.g. “Avoid using vague or general terms”).
  5. 5.
    Grasp: We found that most VLMs other than molmo-7B performed poorly at recognizing keypoints, and instead relied heavily on task information. For example, one rule generated by llava-13b was “A robot should grasp a keypoint that is visible and not obstructed by other objects.” We believe the superior performance of the molmo series of models are due to specifically training on the PixMo dataset with 2D-points (Allen Institute for AI, ).